The perimeter of a square field is 256 yards. How long is each side?

Answer by nyc_function(2741)

Let s = side of your square
Let A = area of your square
The formula for finding the area of a square is A = (side)^2.
A = s^2
256 = s^2
sqrt[256] = sqrt[s^2]
16 = s
Each side is 16 yards long.